OpenStudy (anonymous):

What is the sum 7/2+(-3/5) A. 4/3 b. 29/10 c. -21/10 D/ 41/10

OpenStudy (hba):

Lcm @Lila.hasipi

OpenStudy (aravindg):

did i see this before?

OpenStudy (hba):

Yeah it is the same @AravindG @Lila.hasipi Should i show you an example

OpenStudy (anonymous):

No... @AravindG

OpenStudy (aravindg):

y are u nt able to do it?? i had given u the formula of taking lcm !!

OpenStudy (anonymous):

I think its A

OpenStudy (hba):

\[\frac{ 1 }{ 2 }+(\frac{ -3 }{ 5 })=\frac{ 1 }{ 2 }-\frac{ 3 }{ 5 }\]

OpenStudy (aravindg):

you are right @Lila.hasipi

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

@blurbendy it is!!

OpenStudy (blurbendy):

Start by finding a common denominator for 7/2 and -3/5

OpenStudy (hba):

@AravindG It is not A

OpenStudy (hba):

It is B

OpenStudy (aravindg):

gosh i gt 29/10 and was thinking it was A sorry

OpenStudy (aravindg):

thanks for pointing @blurbendy and @hba

OpenStudy (blurbendy):

:)

OpenStudy (aravindg):

@Lila.hasipi we are not here to do ur homework ..we can assist u ,guide u bt posting worksheets and all is not appreciated